# MIRACLE II: Unveiling the multi-phase gas interplay in the circumnuclear region of NGC 1365 via multi-cloud modeling

### v2

@@ 3.4 Electron density estimation
One of the main advantages of MIRI is the availability of mid-IR density-sensitive line ratios, such as [Ne V]/[Ne V]. These ratios are sensitive to a higher density regime with respect to the commonly used optical [S II]/[S II] ratio, due to the different ionization energies and critical densities of the lines involved. Recent studies used IR tracers to explore the gas density in local Seyfert galaxies and found densities in the range log(ne/cm-3) = 3–5 (Hermosa Muñoz et al., 2024b, a; Zhang et al., 2024; Ceci et al., 2025; Ramos Almeida et al., 2025).
@@
Additionally, as shown in Appendix D, we computed the electron density map from the [S II] doublet observed in the MUSE data (see Venturi et al., 2018 and Mingozzi et al., 2019 for the maps in the entire MUSE FoV). Both the circumnuclear ring and the outflow region are characterized by peaks in electron density, with values reaching log(/cm-3)2.6–2.8. This spatial distribution is consistent with the findings of Venturi et al. (2018) and Kakkad et al. (2018), who reported enhanced electron densities in similar structures. For a quantitative comparison between the optical and mid-IR tracers, we measured the electron density from integrated spectra extracted within a 1.5″ (126 pc) radius aperture centered on the outflow region (see Fig. 6), using both the mid-IR and optical tracers. From the mid-IR [Ne V] line ratio, we estimate an electron density of (750440) cm-3, while from the [S II] lines we find a consistent value of (35040) cm-3. These values are illustrated in the bottom panel of Fig. 6, where we plot the theoretical [S II] and [Ne V] line ratios as functions of electron density using PyNeb (Luridiana et al., 2015), assuming an electron temperature K.
@@ 3.5.1 Model grids of AGN and SF
The other model parameters are identical for the two suites of models (AGN and SF), namely the ionization parameter U (i.e., the number of ionizing photons compared to that of atoms of neutral hydrogen) ranges from to in step of 0.5 and the hydrogen density of the gas, log(NH / cm-3) from 0 to 7 in steps of 1. For the first iteration HOMERUN spans metallicities log(Z/Z⊙) from -1.0 to 0.4 in steps of 0.2. After this first iteration, HOMERUN performs a final refinement in steps 0.02 on predictions interpolated across the finer grid.
@@ 3.5.2 HOMERUN modelling approach
As explained in detail in Marconi et al. (2024), the fitting process starts by selecting a grid of CLOUDY models, defined by hydrogen density () and ionization parameter () for a fixed ionizing spectrum (), gas phase oxygen abundance (), and elemental abundance ratios (). For each such grid, the code searches for the best-fitting multi-cloud model by combining individual single-cloud templates. This is achieved by solving a Non-Negative Least Squares (NNLS) problem, where the weights of each cloud are constrained to be non-negative. The procedure is conceptually analogous to stellar population fitting methods (e.g., pPXF; Cappellari & Emsellem 2004), where a galaxy spectrum is reproduced as a linear combination of stellar templates with positive weights. No regularization is applied in the minimization, allowing for full flexibility in the weight distribution. The process is repeated across all available grids, spanning different ionizing continua, metallicities, and abundance scalings. The optimal solution is identified as the one minimizing the loss function , which quantifies the deviation between model predictions and observed emission-line fluxes. This loss function is effectively a reduced statistic, so values indicate a statistically good fit.
When exploring different metallicities, elemental abundances were scaled from the solar photospheric values of Asplund et al. (2021), except for carbon and nitrogen, which were rescaled following the prescriptions of Nicholls et al. (2017). During the fit, all elemental abundances except oxygen were allowed to vary from their initial values. Scaling factors were used for the AGN and SF components, which could be different for refractory elements but the same for non-refractory ones.
@@ 3.5.3 HOMERUN fitting results
We fit the fluxes of 60 ionized and neutral atomic emission lines measured from the MIRI and MUSE spectra (32 from MIRI and 28 from MUSE, see Figs. 2 and 9) extracted within a 1.5″ radius aperture centered on the outflow region (see Fig. 6). The unique approach of HOMERUN enables us to reproduce at the same time 60 different emission lines in different ionisation stages, with the majority of theme reproduced with better than 15 accuracy. The data are best matched by a SF component with mean = 3.3, / cm-3) = 1.3, and = -3.3 and an AGN component with mean = 1.4, / cm-3) = 2.9, and = -1.4 (see Fig. 7). Being results of ionized gas region, in the following we will consider .
@@ 3.5.4 Ionized gas mass from HOMERUN
After applying HOMERUN to NGC 1365 emission line fluxes, we found the following ionized mass:
@@ 3.6 Outflow energetics
We compare the outflow mass obtained with HOMERUN modeling in the previous section with that estimated following a standard and commonly used approach (e.g. Cano-Díaz et al. 2012; Carniani et al. 2015, see Appendix G). This classical method relies on several strong assumptions, such as considering that all the species are in the ionization state responsible for the observed line (e.g. O2+ for [O III], Ne4+ for [Ne V]), adopting a single average density (from [S II] in the optical and [Ne V] in the mid-IR), and correcting for attenuation using the Balmer decrement.
With HOMERUN we obtain an outflow mass of = (109 2) 103 and = (88 9) 103 within a 1.5″ radius aperture centered on the outflow region (see Fig. 6 and Section 3.5.3). In addition, following Carniani et al. (2015), in Appendix G we present a detailed derivation of the outflow mass of [Ne V] and [O III]. Overall, with the classic method we find in the same spatial region = (2.6 1.0) 103 and = (20 7) 104 . The large difference between the HOMERUN and classical mass estimates (factors of 40 for [Ne V] and 2 for [O III]) clearly highlights the fundamental limitations of the standard approach. The classical method relies on several strong assumptions, such as considering that all the species are in the ionization state responsible for the observed line (e.g. O2+ for [O III], Ne4+ for [Ne V]), adopting a single average density (from [S II] in the optical and [Ne V] in the mid-IR), and correcting for attenuation using the Balmer decrement. In contrast, HOMERUN does not require assumptions about ionization levels, temperature, or metallicity, and accounts for differential reddening between the two components. The reliability of the more complex HOMERUN approach is demonstrated by the fact that classical method mass estimates differ by a factor of 75 between different tracers, highlighting the inherent inconsistencies of the traditional approach. Indeed, for example, the HOMERUN fitting shows that the fraction of neon in the form of Ne4+ in the best-fit models is only 10, while for O2+ the fraction is significantly higher, around 60 of the total oxygen. This is certainly one of the main factors contributing to the large discrepancy between the two mass estimates.
@@
We infer the [Ne V] and [O III] intrinsic (i.e. de-projected) radial velocities of the outflow (vout) using the 3D multi-cloud kinematic model (Marconcini et al., 2023). This model has demonstrated exceptional power in reproducing outflow features with unprecedented detail, leveraging IFU data from a variety of instruments (Marconcini et al., 2023; Cresci et al., 2023; Perna et al., 2024; Ulivi et al., 2025; Marconcini et al., 2025a, b). To reproduce the outflow features in NGC 1365, we adopted a conical outflow morphology with maximum extension of 5 arcsec (420 pc) and position angle111The position angle is measured clockwise from North. of 235∘. The free parameters of the fit are the inclination of the cone axis with respect to the line of sight () and the intrinsic outflow velocity (vout). With we reproduce the line emission on a spaxel-by-spaxel basis, which as a consequence guarantee that the model reproduces the observed moment maps, with discrepancies between observed and modeled moment maps below 5. We found that the best-fit parameters that reproduce the [Ne V] outflow features in NGC 1365 are = 82 6∘, and vout = 390 35 km s-1, consistent with the results in Marconcini et al. (2025b). Fig. 8 shows the 3D representation of this conical outflow. Moreover, we used to fit the [O III] spatially resolved emission assuming the same outflow geometry as for the [Ne V] emission and found = 84 5∘, and vout = 470 40 km s-1. These results are also valid in the region where we extracted the spectrum used for the HOMERUN fitting.
So, adopting R = 252 pc ( 3″, i.e. the sky-projected extension of the 1.5″ radius aperture), we obtain the results of the outflow energetics reported in Table 1. From HOMERUN, we calculate consistent mass outflow rates for [Ne V] and [O III], namely = M⊙ yr-1 and = M⊙ yr-1, respectively. Instead, with the classic approach the [O III]-derived outflow rate is two orders of magnitude larger than that from [Ne V]. This discrepancy is primarily driven by the underestimation of the [Ne V] outflow mass in the classical approach. Specifically, the classical [Ne V] outflow mass is 75 times smaller than the [O III] estimate and 40 times smaller than the HOMERUN value, while the classical [O III] mass is only a factor of 2 higher than the HOMERUN mass. This asymmetric bias clearly demonstrates that the limitations of the standard method predominantly affects high-ionization tracers like [Ne V], leading to severely underestimated masses and consequently unreliable mass outflow rate calculations.
@@ 4 Conclusions
In this work, we presented JWST/MIRI MRS observations of the active galaxy NGC 1365 as part of the Mid-IR Activity of Circumnuclear Line Emission (MIRACLE) program. We analyzed the spatially resolved gas properties and compared them with optical and mm observations from MUSE and ALMA, respectively. We traced both the ionized atomic and the warm and cold molecular gas phases in the circumnuclear region (1.1 1.0 kpc2) of the NGC 1365 galaxy, identifying more than 40 mid-IR emission lines. Our main results are the following:
@@
The ionized gas lines can be separated in two main groups which show opposed morphologies based on their ionization potential (IP). The high-IP (80 V) emission lines trace the bipolar outflow, showing a velocity gradient perpendicular to the disk major axis, consistent with the [O III] kinematics. The velocity dispersion map highlights a donut-shaped structure associated with the approaching side of the outflow which extends up to 2.5″ ( 200 pc). The low-IP (25 eV) emission line kinematics resembles the stellar galaxy disk ordered motions, similar to the H2, with clumpy flux peaks in the circumnuclear star-forming ring. Finally, medium-IP lines show intermediate properties, with a kinematics similar to low-IP species but exhibiting the donut-shaped structure typical of the high-IP lines in the velocity dispersion map.
@@ Acknowledgements.
- Hermosa Muñoz et al. (2024b) Hermosa Muñoz, L., Alonso-Herrero, A., Labiano, A., et al. 2024b, arXiv e-prints, arXiv:2412.14707
@@
- Marconcini et al. (2025a) Marconcini, C., Feltre, A., Lamperti, I., et al. 2025a, arXiv e-prints, arXiv:2503.21921
- Marconcini et al. (2025b) Marconcini, C., Marconi, A., Cresci, G., et al. 2025b, Nature Astronomy [arXiv:2503.24359]
@@ A.1 JWST/MIRI
We refer to Marconcini et al. (2025a) for a detailed description of the different steps performed during the data reduction and briefly summarize them here. We downloaded the uncalibrated science and background observations through the Barbara A. Mikulski Archive for Space Telescopes (MAST) portal333https://www.vla.nrao.edu/astro/nvas/avla.shtml. The data reduction process was performed using the JWST Science Calibration Pipeline (Bushouse et al. 2022) version with JWST Calibration Reference Data System context file 1298. We applied all three stages of the pipeline processing, which include CALWEBB_DETECTOR1, CALWEBB_SPEC2, and CALWEBB_SPEC3 (see Argyriou et al. 2020, 2023; Morrison et al. 2023; Patapis et al. 2024). For NGC 1365, we subtracted the background emission from the 2D science images using a background frame generated from our dedicated background observations. Moreover, we performed two residual fringe corrections, both in stage 2 and 3. Additionally, to improve the final datacubes, we adopted the Exponential Modified-Shepard Method (EMSM) weighting function, which has proven to be more efficient in reducing the drizzling effect (Law et al. 2023).
@@ A.2 MUSE Wide Field Mode
The data were retrieved from the ESO archive444https://archive.eso.org/cms.html and were already processed automatically using the standard MUSE pipeline (v1.6), with an average PSF FWHM of 0.8″. The final datacube covers a FoV of 64″ 63.4″, with a pixel size of 0.2″ pixel-1. The MUSE FoV covers approximately the central 5.3 5.3 kpc2 of NGC 1365. MUSE resolving power in Wide Field Mode (WFM) varies from 1770 at 4800 Å to 3590 at 9300 Å. Finally, we apply a relative astrometric correction, registering the position of the continuum peak, measured in the line-free 7660-8000 Å wavelength range, to the nucleus position from the ALMA data (see Appendix A.3), as done for JWST/MIRI data (see Appendix A.1). The applied corrections were RA and Dec shifts of (-0.78″, +0.22″).

### v3

@@ 3.4 Electron density estimation
One of the main advantages of MIRI is the availability of mid-IR density-sensitive line ratios, such as [Ne V]/[Ne V]. These ratios are sensitive to a higher density regime with respect to the commonly used optical [S II]/[S II] ratio, due to the different ionization energies and critical densities of the lines involved. Recent studies used IR tracers to explore the gas density in local Seyfert galaxies and found densities in the range log(ne/cm-3) = 3–5 (Hermosa Muñoz et al., 2024a, b; Zhang et al., 2024; Ceci et al., 2025; Ramos Almeida et al., 2025).
@@
Additionally, as shown in Appendix D, we computed the electron density map from the [S II] doublet observed in the MUSE data (see Venturi et al., 2018 and Mingozzi et al., 2019 for the maps in the entire MUSE FoV). Both the circumnuclear ring and the outflow region are characterized by peaks in electron density, with values reaching log(/cm-3)2.6–2.8. This spatial distribution is consistent with the findings of Venturi et al. (2018) and Kakkad et al. (2018), who reported enhanced electron densities in similar structures. For a quantitative comparison between the optical and mid-IR tracers, we measured the electron density from integrated spectra extracted within a 1.5″ ( 140 pc) radius aperture centered on the outflow region (see Fig. 6), using both the mid-IR and optical tracers. From the mid-IR [Ne V] line ratio, we estimate an electron density of (750440) cm-3, while from the [S II] lines we find a consistent value of (35040) cm-3. These values are illustrated in the bottom panel of Fig. 6, where we plot the theoretical [S II] and [Ne V] line ratios as functions of electron density using PyNeb (Luridiana et al., 2015), assuming an electron temperature K.
@@ 3.5.1 Model grids of AGN and SF
The other model parameters are identical for the two suites of models (AGN and SF), namely the ionization parameter log(U) (i.e., the number of ionizing photons compared to that of atoms of neutral hydrogen) ranges from to in step of 0.5 and the hydrogen density of the gas, log(nH / cm-3) from 0 to 7 in steps of 1. For the first iteration HOMERUN spans metallicities log(Z/Z⊙) from -1.0 to 0.4 in steps of 0.2. After this first iteration, HOMERUN performs a final refinement in steps 0.02 on predictions interpolated across the finer grid.
@@ 3.5.2 HOMERUN modelling approach
As explained in detail in Marconi et al. (2024), the fitting process starts by selecting a grid of CLOUDY models, defined by hydrogen density () and ionization parameter () for a fixed ionizing spectrum (), gas phase oxygen abundance (), and elemental abundance ratios (). For each grid, the code searches for the best-fitting multi-cloud model by combining individual single-cloud templates. This is achieved by solving a Non-Negative Least Squares (NNLS) problem, where the weights of each cloud are constrained to be non-negative. The procedure is conceptually analogous to stellar population fitting methods (e.g., pPXF; Cappellari & Emsellem 2004), where a galaxy spectrum is reproduced as a linear combination of stellar templates with positive weights. No regularization is applied in the minimization, allowing for full flexibility in the weight distribution. The process is repeated across all available grids, spanning different ionizing continua, metallicities, and abundance scalings. The optimal solution is identified as the one minimizing the loss function , which quantifies the deviation between model predictions and observed emission-line fluxes. This loss function is effectively a reduced statistic, so values indicate a statistically good fit.
When exploring different metallicities, elemental abundances were scaled from the solar photospheric values of Asplund et al. (2021), except for carbon and nitrogen, which were rescaled following the prescriptions of Nicholls et al. (2017). During the fit, all elemental abundances except oxygen were allowed to vary from their initial values. Scaling factors were used for the AGN and SF components. Since the AGN component was assumed to be dust-free while the SF component includes dust, refractory elements can have different scaling factors (due to depletion), whereas non-refractory elements share the same scaling factor in both components.
@@ 3.5.3 HOMERUN fitting results
We fit the fluxes of 60 ionized and neutral atomic emission lines measured from the MIRI and MUSE spectra (32 from MIRI and 28 from MUSE, see Figs. 2 and 9) extracted within a 1.5″ radius aperture centered on the outflow region (see Fig. 6). This aperture was chosen as the largest circular region that is fully covered by all MIRI channels, ensuring the detection of the maximum number of emission lines. The unique approach of HOMERUN enables us to reproduce at the same time 60 different emission lines in different ionisation stages, with the majority of them reproduced with better than 15 accuracy. The data are best matched by a SF component with mean = 3.3, / cm-3) = 1.3, and = -3.3 and an AGN component with mean = 1.4, / cm-3) = 2.9, and = -1.4 (see Fig. 7). Being results of ionized gas region, in the following we will consider .
@@
Using the same procedure, we fitted with HOMERUN the emission lines in the spectra extracted from the regions marked in Fig. 1 (see Figs. 2 and 9). In the outflow-dominated region (green spectrum), we find for the SF component = 2.8, / cm-3) = 0.9, and = -3.4, and for the AGN component = 1.4, / cm-3) = 3.3, and = -1.4. In the stellar disk region (blue spectrum), we find for the SF component = 2.9, / cm-3) = 1.4, and = -3.3, and for the AGN component = 1.5, / cm-3) = 5.5, and = -1.5. These values are consistent with those obtained from the spectrum extracted within a 1.5″ radius aperture (see Fig. 6). Moreover, SF emission accounts for 64 of the H luminosity in the outflow region and 88 in the stellar disk region.
@@ 3.5.4 Ionized gas mass from HOMERUN
After applying HOMERUN to NGC 1365 emission line fluxes, we found the following ionized gas masses:
@@ 3.6 Outflow energetics
In the previous section, with HOMERUN we obtain an outflow mass of = (109 2) 103 and = (88 9) 103 within a 1.5″ radius aperture centered on the outflow region (see Fig. 6 and Section 3.5.3). In addition, following Carniani et al. (2015), in Appendix G we present a detailed derivation of the outflow mass of [Ne V] and [O III]. Overall, with the classic method we find in the same spatial region = (2.6 1.0) 103 and = (20 7) 104 . The large difference between the HOMERUN and classical mass estimates (factors of 40 for [Ne V] and 2 for [O III]) clearly highlights the fundamental limitations of the standard approach. The classical method relies on several strong assumptions, such as considering that all the species are in the ionization state responsible for the observed line (e.g. O2+ for [O III], Ne4+ for [Ne V]), adopting a single average density (from [S II] in the optical and [Ne V] in the mid-IR), and correcting for attenuation using the Balmer decrement. In contrast, HOMERUN does not require assumptions about ionization levels, temperature, or metallicity, and accounts for differential reddening between the two components. The reliability of the more complex HOMERUN approach is demonstrated by the fact that classical method mass estimates differ by a factor of 75 between different tracers, highlighting the inherent inconsistencies of the traditional approach. Indeed, for example, the HOMERUN fitting shows that the fraction of neon in the form of Ne4+ in the best-fit models is only 10, while for O2+ the fraction is significantly higher, around 60 of the total oxygen. This is certainly one of the main factors contributing to the large discrepancy between the two mass estimates. Moreover, it should be noted that [O III] and [Ne V], having different critical densities and luminosities, trace gas from slightly different physical regions within the outflowing clouds, which further contributes to the observed differences in mass estimates beyond the methodological inconsistencies of the classical approach.
@@
We infer the [Ne V] and [O III] intrinsic (i.e. de-projected) radial velocities of the outflow (vout) using the 3D multi-cloud kinematic model (Marconcini et al., 2023). This model has demonstrated exceptional power in reproducing outflow features with unprecedented detail, leveraging IFU data from a variety of instruments (Marconcini et al., 2023; Cresci et al., 2023; Perna et al., 2024; Ulivi et al., 2025; Marconcini et al., 2025a, b). To reproduce the outflow features in NGC 1365, we adopted a conical outflow morphology with maximum extension of 5 arcsec ( 470 pc) and position angle333The position angle is measured clockwise from North. of 235∘. The free parameters of the fit are the inclination of the cone axis with respect to the line of sight () and the intrinsic outflow velocity (vout). With we reproduce the line emission on a spaxel-by-spaxel basis, which as a consequence guarantee that the model reproduces the observed moment maps, with discrepancies between observed and modeled moment maps below 5. We found that the best-fit parameters that reproduce the [Ne V] outflow features in NGC 1365 are = 82 6∘, and vout = 390 35 km s-1, consistent with the results in Marconcini et al. (2025b). Fig. 8 shows the 3D representation of this conical outflow. Moreover, we used to fit the [O III] spatially resolved emission assuming the same outflow geometry as for the [Ne V] emission and found = 84 5∘, and vout = 470 40 km s-1. These results are also valid in the region where we extracted the spectrum used for the HOMERUN fitting.
So, adopting R = 284 pc ( 3″, i.e. the sky-projected extension of the 1.5″ radius aperture), we obtain the results of the outflow energetics reported in Table 1. From HOMERUN, we calculate consistent mass outflow rates for [Ne V] and [O III], namely = M⊙ yr-1 and = M⊙ yr-1, respectively. Instead, with the classic approach the [O III]-derived outflow rate is two orders of magnitude larger than that from [Ne V]. This discrepancy is primarily driven by the underestimation of the [Ne V] outflow mass in the classical approach. Specifically, the classical [Ne V] outflow mass is 75 times smaller than the [O III] estimate and 40 times smaller than the HOMERUN value, while the classical [O III] mass is only a factor of 2 higher than the HOMERUN mass. This asymmetric bias clearly demonstrates that the limitations of the standard method predominantly affects high-ionization tracers like [Ne V], leading to severely underestimated masses and consequently unreliable mass outflow rate calculations.
@@ 4 Conclusions
In this work, we presented JWST/MIRI MRS observations of the active galaxy NGC 1365 as part of the Mid-IR Activity of Circumnuclear Line Emission (MIRACLE) program. We analyzed the spatially resolved gas properties and compared them with optical and mm observations from MUSE and ALMA, respectively. We traced both the ionized atomic and the warm and cold molecular gas phases in the circumnuclear region ( 0.90.9 kpc2) of the NGC 1365 galaxy, identifying more than 40 mid-IR emission lines. Our main results are the following:
@@
The ionized gas lines can be separated in two main groups which show opposed morphologies based on their ionization potential (IP). The high-IP ( 54 eV) emission lines trace the bipolar outflow, showing a velocity gradient perpendicular to the disk major axis, consistent with the [O III] kinematics. The velocity dispersion map highlights a donut-shaped structure associated with the approaching side of the outflow which extends up to 2.5″ ( 240 pc). The low-IP ( 25 eV) emission line kinematics resembles the stellar galaxy disk ordered motions, similar to the H2, with clumpy flux peaks in the circumnuclear star-forming ring. Finally, medium-IP lines show intermediate properties, with a kinematics similar to low-IP species but exhibiting the donut-shaped structure typical of the high-IP lines in the velocity dispersion map.
@@ Acknowledgements.
- Fitzpatrick et al. (2019) Fitzpatrick, E. L., Massa, D., Gordon, K. D., Bohlin, R., & Clayton, G. C. 2019, ApJ, 886, 108
@@
- Gordon (2024) Gordon, K. 2024, The Journal of Open Source Software, 9, 7023
- Gordon et al. (2009) Gordon, K. D., Cartledge, S., & Clayton, G. C. 2009, ApJ, 705, 1320
@@
- Hermosa Muñoz et al. (2024b) Hermosa Muñoz, L., Alonso-Herrero, A., Pereira-Santaella, M., et al. 2024b, A&A, 690, A350
@@
- Marconcini et al. (2025a) Marconcini, C., Feltre, A., Lamperti, I., et al. 2025a, A&A, 701, A113
- Marconcini et al. (2025b) Marconcini, C., Marconi, A., Cresci, G., et al. 2025b, Nature Astronomy, 9, 907
@@ A.1 JWST/MIRI
We refer to Marconcini et al. (2025a) for a detailed description of the different steps performed during the data reduction and briefly summarize them here. We downloaded the uncalibrated science and background observations through the Barbara A. Mikulski Archive for Space Telescopes (MAST) portal444https://www.vla.nrao.edu/astro/nvas/avla.shtml. The data reduction process was performed using the JWST Science Calibration Pipeline (Bushouse et al. 2022) version with JWST Calibration Reference Data System context file 1298. We applied all three stages of the pipeline processing, which include CALWEBB_DETECTOR1, CALWEBB_SPEC2, and CALWEBB_SPEC3 (see Argyriou et al. 2020, 2023; Morrison et al. 2023; Patapis et al. 2024). For NGC 1365, we subtracted the background emission from the 2D science images using a background frame generated from our dedicated background observations. Moreover, we performed two residual fringe corrections, both in stage 2 and 3. Additionally, to improve the final datacubes, we adopted the Exponential Modified-Shepard Method (EMSM) weighting function, which has proven to be more efficient in reducing the drizzling effect (Law et al. 2023).
@@ A.2 MUSE Wide Field Mode
The data were retrieved from the ESO archive555https://archive.eso.org/cms.html and were already processed automatically using the standard MUSE pipeline (v1.6), with an average PSF FWHM of 0.8″. The final datacube covers a FoV of 64″ 63.4″, with a pixel size of 0.2″ pixel-1. The MUSE FoV covers approximately the central 5.3 5.3 kpc2 of NGC 1365. MUSE resolving power in Wide Field Mode (WFM) varies from 1770 at 4800 Å to 3590 at 9300 Å. Finally, we apply a relative astrometric correction, registering the position of the continuum peak, measured in the line-free 7660-8000 Å wavelength range, to the nucleus position from the ALMA data (see Appendix A.3), as done for JWST/MIRI data (see Appendix A.1). The applied corrections were RA and Dec shifts of (-0.78″, +0.22″).